Present Value of an Annuity

Determine the present value of $130,000 to be received at the end of each of four years, using an interest rate of 7%, compounded annually, as follows:

a. By successive computations, using the present value of $1 table in Exhibit 5. Round to the nearest whole dollar.

First year $
Second Year $
Third Year $
Fourth Year $
Total present value $

b. By using the present value of an annuity of $1 table in Exhibit 7. Round to the nearest whole dollar.
$

c. Why is the present value of the four $130,000 cash receipts less than the $520,000 to be received in the future?
The present value is less due to (inflation, the compounding of interest, deflation) over the 4 years.
Answer #1
a) First year $ 1,21,495
Second Year $ 1,13,547
Third Year $ 1,06,119
Fourth Year $     99,176
Total present value $ 4,40,337
Note: Interest factor = 1/1.07 = 0.9346
b) PVIFA(7,4) = 3.3872
Total present value = 130000*3.3872) $    4,40,337
c) The PV is less due to compounding of interest.